OpenStudy (joftheworld):

the length and width of a rectangle are consecutive even integers. the area of the rectangle is 120 square units. what are the length and width ?

OpenStudy (phi):

you need to know a few things: 1) what is the formula for the area of a rectangle?

OpenStudy (joftheworld):

a=wl but i have to solve it with quadratic equations...

OpenStudy (phi):

2) what does "consecutive even integers" mean ?

OpenStudy (joftheworld):

they have to be even ?

OpenStudy (phi):

yes, but more than that. they are "next to" each other. examples: 2 and 4 (we skip the 3 which is odd) or 6 and 8 and so on

OpenStudy (phi):

We don't know what the consecutive integers are, so we have to pick a variable name to represent them. We can do it two ways: pick x and x+2 to stand for the two consecutive integers. (of course, this could stand for two consecutive odd numbers... but if we solve the problem we will probably find two even numbers... it's a made up problem and they know we will get even numbers) or we can force the numbers to be even by using 2x and 2x+2 2x will be even and so will 2x+2

OpenStudy (phi):

if we use 2x and 2x+2 the problem will be slightly easier to solve. So lets do that. we don't know which is the length and which the width, but we can just randomly choose: w= 2x and L= 2x+2 use those definitions in your formula for the area. Can you do that ? Can you write down the formula?

OpenStudy (joftheworld):

2x(2x+2)=120 ?

OpenStudy (phi):

yes. now distribute the 2x. what do you get ?

OpenStudy (joftheworld):

4x^2+2x=120 ?

OpenStudy (phi):

now put it in "standard form" (all terms on the left side)

OpenStudy (joftheworld):

4x^2+4x-120=0

OpenStudy (joftheworld):

sorry i messed up on the other equation lol

OpenStudy (phi):

although you should check your "distribute the 2x step" oh... you have it correct!

OpenStudy (phi):

now it's always good if you can factor a constant out of each term (if you can) to simplify the quadratic. In this case, divide all terms by 4. and to keep things equal, divide the other side (the 0) by 4

OpenStudy (joftheworld):

\[4x ^{2}+4x-120=4*(x-5)*(x+6)\]

OpenStudy (phi):

and 4(x−5)(x+6) =0 if we simplify (divide both sides by 4) (x−5)(x+6) =0 and now we have either (x-5)=0 or x+6= 0 solve for x then use w= 2x and L= 2x+2

OpenStudy (joftheworld):

x=5 x=-6 ?

OpenStudy (phi):

yes, and w= 2x and L= 2x+2

OpenStudy (joftheworld):

so i just simplify that ? or leave it as it is ?

OpenStudy (phi):

we want numbers for the width and length

OpenStudy (joftheworld):

2,4 ?

OpenStudy (phi):

we (or rather, you) found x=5 x=-6 use those values in w= 2x and L= 2x+2

OpenStudy (joftheworld):

i dont get it D:

OpenStudy (phi):

yes, we'll review in a minute. in the meantime, us x=5 in w= 2x and L= 2x+2

OpenStudy (joftheworld):

lol 12 and 10

OpenStudy (anonymous):

@phi ,@joftheworld see the link

OpenStudy (joftheworld):

OH ! i get is now ! lol

OpenStudy (joftheworld):

it*

OpenStudy (phi):

and just to be complete, what do we get for w and L when x= -6?

OpenStudy (joftheworld):

just multiply 5 and -6 by 2 ?

OpenStudy (joftheworld):

but wouldnt it be 10, -12 ?

OpenStudy (phi):

replace x with -6 in w= 2x and L= 2x+2

OpenStudy (joftheworld):

10,10 ?

OpenStudy (joftheworld):

or 10,-10 ?

OpenStudy (phi):

slow down.... just erase the x, put in (times) -6 in w= 2x and L= 2x+2

OpenStudy (joftheworld):

w=2(5) >>> 10 L=2(-6)+2 >>>-10

OpenStudy (phi):

oh. no, we don't use different x's to find w and L we use the *same* x to get w and L we already used x=5 to find w= 2*5 =10 and L= 2*5+2 = 12 now we are seeing what we get for w and L when we use x=-6

OpenStudy (joftheworld):

-12, -10 ?

OpenStudy (phi):

yes, and notice with w= -12 and L=-10 w*L= -12* -10= +120 so this "answer" works... but we don't use negative numbers for lengths... we ignore this "answer" (sometimes called an "extraneous solution" ) the other pair, w=10 and L=12 is the answer. (though we don't know if it's w=12 and L=10 ... though I think of the width as being shorter than the length in a rectangle)

OpenStudy (joftheworld):

thanks can you help with couple more ? it should be easy..

OpenStudy (joftheworld):

the length of a rectangle plus its width is 18 cm, the area is 80 square cm. what is the length and width ?

OpenStudy (joftheworld):

im guessing its 10,8 ? because 80/8=10 ? and 10+8 is 18 ..

OpenStudy (phi):

Now to review: area of the rectangle is 120 we used that to write w*L= 120 length and width of a rectangle are consecutive even integers this part is a bit tricky. To get an *even* number, we write 2x 2x will always be even, no matter what x is. (this is a "trick" people learn) if we have 2x, the next integer up is 2x+2 (btw, we could have gone down, and picked 2x-2... that will work) I said let w be the first number, w= 2x (going on the theory that the width should be smaller than the length.. but it does not have to be) and let L be the other number, L= 2x+2 we then said w*L= 120 2x(2x+2) =120 from which we get 4x^2 +4x -120=0 or x^2 +x -30=0 (x+6)(x-5) =0 x= 5 or x=-6 we use x=5 in w=2x and L=2x+2 (notice we use the same x, or we would not get consecutive even numbers for w and L) we get w=10 and L=12

OpenStudy (joftheworld):

did you read what i said above ?

OpenStudy (phi):

I just did. yes you have the answer. Now can you solve it using equations?

OpenStudy (joftheworld):

yes but i have to hurry up and do this lesson .. i have to be somewhere soon

OpenStudy (phi):

length of a rectangle plus its width is 18 cm write down the formula this is saying

OpenStudy (joftheworld):

10*8=80

OpenStudy (joftheworld):

whoops sorry 10+8=18

OpenStudy (phi):

yes, if you know the answer, you can do that. but if you don't know the numbers, how do you write it?

OpenStudy (joftheworld):

x(x+1)=80 ?

OpenStudy (joftheworld):

are you good with functional values ?

OpenStudy (phi):

if you don't have numbers, you use names (letters) instead in fact, it's the same thing as length of a rectangle plus its width is 18 cm but "shortened" by replacing the long words with short names.